OpenStudy (anonymous):

2 what is the product of ( - 6 ) ? ( 1point ) A . -3 b. -12 c. 36 d. 6

OpenStudy (kirbykirby):

\(a^2=a\times a\) here, \(a=-6\) so: \((-6)^2=(-6)\times(-6)\)